# A comprehensive evaluation of spatial co-execution on GPUs using MPS and MIG technologies

### 24 Apr 2026 v1

@@ 5.1 Energy consumption
In contrast, spatial sharing inevitably increases instantaneous power consumption due to higher resource concurrency. However, this overhead is often cost-effective, as the reduction in execution time frequently offsets the power penalty, leading to overall energy savings (up to 20% in some cases). As shown in Figure 12, MPS exhibits a modest power increase ($\simeq 7\%$) for non-contentious pairs such as SRAD-LavaMD ($P_{\text{rel}}\!=\!1.07$, $T_{\text{rel}}\!=\!0.68$). Here, the trade-off is clearly advantageous, yielding a net energy reduction of over 20% (situated below the 0.8 isocurve). Conversely, for pairs suffering from contention or saturation, such as Sort-Sort or Enformer-Enformer ($P_{\text{rel}}\!=\!1.15$, $T_{\text{rel}}\!\simeq\!1.3$), the power penalty rises to 15%. In these scenarios, the combination of higher power and longer execution time results in an energy degradation exceeding 40% (above the 1.4 isocurve)—a penalty even more severe than the performance loss itself. In conclusion, MPS has the potential to produce highly polarized results, and using it exclusively without resource contention is even more important in terms of energy consumption.
@@
Using MPS with SM provisioning results in a consistently higher power draw versus basic MPS, ranging from 10% to 15% extra. This is potentially attributable to the additional logic required for SM-level isolation. However, there appears to be no significant difference in the overall power-performance trade-off, as the data points for both basic MPS and MPS-best-prov show a similar distribution relative to the isocurves (see particularly SRAD-LavaMD, Sort-Sort and Enformer-Enformer, which are highlighted in both). This suggests that the higher power draw of provisioning is generally offset by corresponding performance gains.
@@
Finally, MIG exhibits the highest power levels in Figure 12: typically around a 15% increase, reaching 25% when handling applications more demanding than the assigned instance such as Enformer and GeMM. This is natural, because full resource isolation must necessarily incur an energy overhead, especially when the instance size is insufficient for the application’s demand as in those cases, causing issues like increased stress on high memory levels (as in the example from Table 4). Although MIG delivered the best average performance (with y-values more frequently below 1.0), its peak performance does not match the best cases of MPS, and conversely, it also avoids severe performance degradation. Consequently, the energy trade-off is many times unfavorable, as approximately half of the MIG test cases fall above the 1.0 isocurve, even though most are below $T_{\text{rel}}=1$ (they improve the original performance). For example, Sort-GUPS, despite reducing execution time by 10% ($T_{\text{rel}}\!=\!0.9$), is not energy-efficient because its power draw increases by 19% ($P_{\text{rel}}\!=\!1.19$). Furthermore, even when MIG generates energy savings, these savings do not reach the levels seen with MPS (no MIG pair falls below the 0.8 isocurve, although LavaMD-LavaMD is close). The conclusion is that MIG hardware isolation carries an important inherent energy cost, and to be energy efficient, the performance gains must be substantial enough to offset this overhead (which occurs in approximately half of our test cases).
@@ 5.2 Performance concurrency scaling
  • Medium workloads: With the same configuration as in the previous tests, stated in Table 1. ResNet50 and BERT will be the RDL models reported because the others have very similar results and these have lightweight versions (ResNet18 and DistilBERT) to evaluate with a higher degree of co-execution.
@@ 5.3 Energy efficiency scaling
Regarding the Energy-Delay Product (EDP), the results in Figure 15 show that performance improvement successfully compensates for the increased energy consumption with four concurrent copies. As shown in Table 5, the reduction in execution time was approximately 20% compared to the sequential baseline, while the corresponding energy increase was only around 10% (see in Figure 14 Sort, GUPS, KMeans and SRAD, especially with MPS, and ResNet18 and DistilBERT with MIG). Consequently, the overall EDP, which gives equal weight to both factors, is favorable ($EDP_{\text{rel}}<1$), indicating that the gains in performance outweigh the energy penalty (for all applications except GeMM). The application-by-application analysis reveals a combination of the phenomena already discussed, and it is striking how GeMM’s EDP is extremely poor with 4 copies (not so with 2, with an enormous difference between them) because its demand is excessive for that degree of concurrency, which impairs performance but also increases consumption (unlike other tests where compensatory trade-offs were observed).
5.4 Dynamic reconfiguration
Until now, MPS and MIG have been exhaustively analyzed from both performance and energy perspectives under various static configurations (initially with co-executed pairs, and subsequently with sets of four applications). However, it remains to evaluate the change of configurations, which can be done “on the fly”, i.e., while some applications are running (dynamic reconfiguration).
In the case of MPS, the inherent cost of such reconfigurations is practically negligible. Whether introducing a new application into an existing MPS context or varying the SM limit percentage (in the provisioned version), no overhead is observed in application launch relative to the baseline cost (i.e., launch latency or power consumption prior to the reconfiguration request). A separate issue is whether the new configuration yields better or worse performance and consumption, a factor already analyzed in our previous study of static configurations, specifically regarding the contention problems and the scaling from 2 to 4 concurrent applications.
Table 6: MIG reconfiguration latency and power consumption overhead (relative to 32W of idle GPU) by instance size and operation type (average of 10 runs per test).
$t_{\text{reconfig.}}$(s) $\vartriangle\!\text{Power}$ (W) $\vartriangle\!\text{Energy}$ (J)
Slices Create Destroy Create Destroy Create Destroy
1 0.11 0.10 21.12 17.54 2.32 1.75
2 0.12 0.10 19.64 17.71 2.36 1.77
4 0.13 0.10 18.95 20.82 2.46 2.08
In the case of MIG, reconfiguration entails a more significant cost because it involves not only a logical modification of resource distribution but also a hardware alteration of the grouping of fully isolated instances (i.e., their physical destruction and creation). Table 6 summarizes the duration of MIG instance destruction and creation according to its size on an NVIDIA A30 ($t_{\text{reconfig.}}$), as well as the increase in power ($\vartriangle\!\text{Power}$) and energy consumption ($\vartriangle\!\text{Energy}$) relative to baseline values (the idle GPU consumes a stable 32W at fixed operating frequencies). As observed, the overhead is quite low, with latencies of merely a few tenths of a second per instance—affecting only the slices of that specific instance and not those of others running applications—and power increases of $\sim\!20$W, which translates to $\sim\!2$J of extra energy per operation. This makes the use of reconfigurations to better adapt resource distribution to workload characteristics highly feasible for performance optimization, as demonstrated in our previous MIG scheduling proposals [VillarrubiaFAR2025, VillarrubiaDRL2026]. Furthermore, the overhead is shown to be practically independent of instance size, as creating or destroying instances of size 1, 2, or 4 incurs nearly identical costs.
Finally, it has been empirically verified that the performance of multiple reconfigurations is equivalent to sequential execution, even when attempted in parallel (where one might a priori expect parallelization over disjoint resources). This means that even if distinct CPU threads or processes attempt to simultaneously create or destroy disjoint instances on the GPU, the total duration is approximately the sum of the individual operations. This behavior has been verified for all possible partition transitions on A30 and A100 GPUs using NVIDIA driver version 580.82.07 (subsequent versions show no changes in this regard). This might be improved in the future, unless there are inherent technical limitations unknown to us.
@@ 7 Conclusions
This paper has presented a comprehensive empirical evaluation of spatial co-execution on modern GPUs, focusing on NVIDIA’s MPS and MIG technologies. We have compared these against a standard time-slicing baseline, analyzing their performance, energy consumption, and scalability across a diverse set of workloads with varying computational and memory demands. Our findings highlight the significant potential of spatial sharing to improve GPU utilization and efficiency, while also revealing the critical trade-offs between the flexibility of MPS and the strong isolation of MIG. GPUs and technologies from other manufacturers such as AMD and Intel work differently and would require their own evaluation, but NVIDIA is more relevant (it is the clear market leader) and the basic principles regarding the trade-offs between sharing-based and isolation-based co-execution are, a priori, extrapolable. The key insights obtained are summarized below.
@@ 7.1 Key insights
    MIG provides robust performance isolation by partitioning hardware resources, which effectively prevents interference between co-running applications, at the cost of a minor overhead during dynamic reconfiguration (Section 5.4). Consequently, most application pairs achieved a notable performance improvement compared to sequential execution, with many showing a 20% reduction in overall slowdown and some reaching up to 30% (see Figure 12). This benefit was particularly stark for combinations of two memory-intensive (M-type) workloads, which completed their execution almost twice as fast as with MPS (see Figure 6).
@@
    MPS offers a more flexible approach to resource sharing, allowing for fine-grained allocation of computational resources. This flexibility, however, comes at the cost of potential contention for shared resources, particularly the memory subsystem. The ability to provision SMs provides a mechanism to mitigate some SM-level sharing issues and, crucially, to prevent resource monopolization, as was observed with the dominant GeMM application in Figure 10 and the Enformer model in Figure 11. Nevertheless, this provisioning does not solve the underlying memory contention problems inherent in memory-intensive (M-type) applications. As a result, the performance and energy consumption outcomes with MPS are highly polarized (see Figure 12). It achieves better results than MIG in favorable scenarios but performs significantly worse when memory conflicts arise. This polarization was further accentuated when scaling to four concurrent applications (Figures 5 and 14). These findings underscore the critical importance of careful workload characterization to select the appropriate technology.
@@
    The evolution of MIG would benefit from more granular and flexible partitioning schemes to reduce resource underutilization. Specifically, the ability to create instances with asymmetric resources (e.g., more memory bandwidth relative to compute resources) would better match diverse application profiles. While technically challenging in the short term, this would help bridge the gap between MIG’s strong isolation and MPS’s flexibility. Similarly, dynamic scheduling would benefit if reconfiguration operations on disjoint slices could be effectively parallelized, given that their current performance is equivalent to sequential reconfiguration (Section 5.4).
@@ 7.2 Future work
J.V. conducted the design, implementation and evaluation of the experiments, and wrote the main manuscript text. L.C., F.I. and K.O. collaborated on the definition and supervision of research tasks, the design and critical analysis of the experiments, and the review and writing of the manuscript.

### 29 Apr 2026 v2

@@ 5.1 Energy consumption
In contrast, spatial sharing methods increase power draw more significantly (which is expected due to their higher use of resources), but in many cases this is cost-effective, as performance improves enough to reduce overall energy consumption (by up to 20% in some cases). In Figure 12, MPS shows power increases of around 7% for pairs without contention or resource monopolization issues, such as SRAD-LavaMD ($P_{\text{rel}}\!=\!1.07$, $T_{\text{rel}}\!=\!0.68$). This increase reaches up to 15% for pairs that do exhibit these problems, like Sort-Sort (an M-M combination with memory contention and $P_{\text{rel}}\!=\!1.15$, $T_{\text{rel}}\!=\!1.31$). These wide performance variations, coupled with the differing power draws, lead to highly variable energy outcomes. This results in cases with over 20% energy improvement, such as SRAD-LavaMD (below the 0.8 isocurve), and others like Sort-Sort with a degradation of over 40% (above the 1.4 isocurve). The conclusion is that co-execution with MPS can yield extreme outcomes, both positive and negative. It is therefore critical to avoid contention scenarios, a conclusion that is even more pronounced for energy consumption than it was for performance.
@@
Using MPS with SM provisioning results in a consistently higher power draw versus basic MPS, ranging from 10% to 15% extra. This is potentially attributable to the additional logic required for SM-level isolation. However, there appears to be no significant difference in the overall power-performance trade-off, as the data points for both basic MPS and MPS-best-prov show a similar distribution relative to the isocurves (see particularly SRAD-LavaMD, which is highlighted in both). This suggests that the higher power draw of provisioning is generally offset by corresponding performance gains.
@@
Finally, MIG exhibits the highest power levels in Figure 12: typically around a 15% increase, reaching 25% when handling applications more demanding than the assigned instance such as GeMM-GeMM. This is natural, because full resource isolation must necessarily incur an energy overhead, especially when the instance size is insufficient for the application’s demand as in the GeMM case and causes issues like increased stress on high memory levels (as in the example from Table 4). Although MIG delivered the best average performance (with y-values more frequently below 1.0), its peak performance does not match the best cases of MPS, and conversely, it also avoids severe performance degradation. Consequently, the energy trade-off is many times unfavorable, as approximately half of the MIG test cases fall above the 1.0 isocurve, even though most are below $T_{\text{rel}}=1$ (they improve the original performance). For example, Sort-GUPS, despite reducing execution time by 10% ($T_{\text{rel}}\!=\!0.9$), is not energy-efficient because its power draw increases by 19% ($P_{\text{rel}}\!=\!1.19$). Furthermore, even in cases where MIG does yield energy savings, these savings do not reach the levels seen with MPS (no MIG pair falls below the 0.8 isocurve, although LavaMD-LavaMD is close). The conclusion is that MIG hardware isolation carries an important inherent energy cost, and to be energy efficient, the performance gains must be substantial enough to offset this overhead (which occurs in approximately half of our test cases).
@@ 5.2 Performance concurrency scaling
  • Medium workloads: Maintaining the same configuration as in the previous tests, stated in Table 1.
@@ 5.3 Energy efficiency scaling
Regarding the Energy-Delay Product (EDP), the results in Figure 15 show that performance improvement successfully compensate for the increased energy consumption with four concurrent copies. As shown in Table 5, the reduction in execution time was approximately 20% compared to the sequential baseline, while the corresponding energy increase was only around 10% (see in Figure 14 Sort, GUPS, KMeans and SRAD, especially with MPS, and ResNet18 and DistilBERT with MIG). Consequently, the overall EDP, which gives equal weight to both factors, is favorable ($EDP_{\text{rel}}<1$), indicating that the gains in performance outweigh the energy penalty (for all application except GeMM). The application-by-application analysis reveals a combination of the phenomena already discussed, and it is striking how GeMM’s EDP is extremely poor with 4 copies (not so with 2, with an enormous difference between them) because its demand is excessive for that degree of concurrency, which impairs performance but also increases consumption (unlike other tests where compensatory trade-offs were observed).
@@ 7 Conclusions
This paper has presented a comprehensive empirical evaluation of spatial co-execution on modern GPUs, focusing on NVIDIA’s MPS and MIG technologies. We have compared these against a standard time-slicing baseline, analyzing their performance, energy consumption, and scalability across a diverse set of workloads with varying computational and memory demands. Our findings highlight the significant potential of spatial sharing to improve GPU utilization and efficiency, while also revealing the critical trade-offs between the flexibility of MPS and the strong isolation of MIG. The key insights obtained are summarized below.
@@ 7.1 Key insights
    MIG provides robust performance isolation by partitioning hardware resources, which effectively prevents interference between co-running applications, albeit with a minor overhead. Consequently, most application pairs achieved a notable performance improvement compared to sequential execution, with many showing a 20% reduction in overall slowdown and some reaching up to 30% (see Figure 12). This benefit was particularly stark for combinations of two memory-intensive (M-type) workloads, which completed their execution almost twice as fast as with MPS (see Figure 6).
@@
    MPS offers a more flexible approach to resource sharing, allowing for fine-grained allocation of computational resources. This flexibility, however, comes at the cost of potential contention for shared resources, particularly the memory subsystem. The ability to provision SMs provides a mechanism to mitigate some SM-level sharing issues and, crucially, to prevent resource monopolization, as was observed with the dominant GeMM application in Figure 10. Nevertheless, this provisioning does not solve the underlying memory contention problems inherent in memory-intensive (M-type) applications. As a result, the performance and energy consumption outcomes with MPS are highly polarized (see Figure 12). It achieves better results than MIG in favorable scenarios but performs significantly worse when memory conflicts arise. This polarization was further accentuated when scaling to four concurrent applications (Figures 5 and 14). These findings underscore the critical importance of careful workload characterization to select the appropriate technology.
@@
    The evolution of MIG would benefit from more granular and flexible partitioning schemes to reduce resource underutilization. Specifically, the ability to create instances with asymmetric resource (e.g., more memory bandwidth relative to compute resources) would better match diverse application profiles. While technically challenging in the short term, this would help bridge the gap between MIG’s strong isolation and MPS’s flexibility.
@@ 7.2 Future work
J.V. conducted the design, implementation and evaluation of the experiments, and wrote the main manuscript text. L.C., F.I and K.O collaborated in the definition and supervision of research tasks, the design and critical analysis of the experiments, and the review and writing of the manuscript.